OpenStudy (anonymous):

if Mr.Weidenheft said that 7/8 of his students showed up for the final and he has 32 total students, how many students were absent?

OpenStudy (anonymous):

i have no idea how to figure this out btw

OpenStudy (anonymous):

Absent=Total-showed up ones , 32-32*7/8 so 32-28=4

OpenStudy (anonymous):

quick question, where'd al these numbers come from?

OpenStudy (anonymous):

7/8 showed up, so 1/8 didn't show up. 1/8*32=4

OpenStudy (anonymous):

ohhhh ok thx!!!

OpenStudy (anonymous):

no of pple who didnt show up,1/8*32=4.

OpenStudy (anonymous):

wait a second... you sure that you multiply? 4 is smaller than 32 so why wouldn't you divide?

OpenStudy (anonymous):

never mind i got it

OpenStudy (anonymous):

total is 8/8 and 7/8 were present.so those absent is 8/8-7/8 which is 1/8.so to find those absent just multiply 1/8 by 32